### v1

@@ I Introduction
In relation to the Cauchy horizon and the phenomenon of mass inflation, a complementary quantum-gravitational resolution is proposed: (i) Interpret the interior wave function as imposing an “annihilation-to-nothing” boundary condition, so that both the event and Cauchy horizons lie in the causal past of the bounce. In this picture, the infinite blueshift at the would-be Cauchy horizon is not operationally relevant. () With only an ingoing mode from the event horizon, no physical structure forms at the Cauchy horizon. While not a classical completion, an appropriate boundary condition on the wave function smoothly ameliorates the interior geometry and removes the mechanism for mass inflation. Whether this mechanism extends to rotating or regular black holes with inner horizons (e.g., Kerr or regular charged solutions) remains open, and we leave a systematic treatment of these cases to future work.
@@ II Wheeler-DeWitt Equation for Static Black Hole Interiors
where $r_{Q}^{2}={q^{2}}/{4\pi\epsilon_{0}}$ is a characteristic length scale with respect to charge $q$. The metric ansatz of the spacetime inside a static black hole takes the form
@@ IV.3 Wave Function and Its Boundary Conditions
We note that the choice of the width parameter $\sigma$ for the wave function $\Psi_{-}$ presents a numerical challenge. A small value of $\sigma$ yields a sharply localized Gaussian but can lead to numerical instability, while a large $\sigma$ improves numerical stability at the cost of producing an undesirably broad wave packet. In the regime $r_{Q}\ll r_{s}$, the inner-horizon radius $r_{-}$ becomes extremely small, so that the Gaussian factor $\exp(\sigma^{2}(\lambda-r_{-})^{2}/2)$ is effectively displaced from the intended location. To mitigate this, we adopt a larger $\sigma$ together with a larger $r_{Q}$. This adjustment, however, reduces the decay rate of the wave amplitude in the QQ–direction, as seen in Fig. 3, leading to partial overlap between the two horizon-sourced components. This overlap should not be interpreted as a genuine physical effect of the Reissner–Nordström interior, but rather as a consequence of numerical limitations; in the ideal physical scenario, the two waves from $r_{\pm}$ would remain sharply localized and well-separated, as depicted in Fig. 2.
@@ IV.4 Neutral Black Hole Limit
In the monotonic‑decay scenario, Schwarzschild and Schwarzschild–(anti)‑de Sitter Wheeler–DeWitt solutions satisfy the DeWitt boundary condition at the curvature singularity, realizing singularity avoidance; however, prior analyses achieve this at the cost of an off‑shell growth that makes the wave function unbounded (see Sec. III.D of Ref. Bouhmadi-Lopez:2019kkt and Sec. III.A of Ref. Chien:2023kqw under the purely incoming horizon condition). In contrast, by reinterpreting the Schwarzschild black hole as the charge‑neutral limit of the Reissner–Nordström black hole and implementing a Gaussian wave with charge fluctuations as in Eq. (33), the neutral‑limit state is normalizable and exhibits strict monotonic decay. This construction furnishes a bounded realization of the monotonic‑decay solution while preserving the DeWitt boundary condition, therefore providing a new perspective on neutralization and refining the late‑time Schwarzschild end‑state picture (cf. Refs. Bouhmadi-Lopez:2019kkt ; Chien:2023kqw ), as illustrated in Fig. 4.
@@ VI Comments on Cauchy Horizon and Mass Inflation
Furthermore, if we think that there is only one piece of the incoming wave from the event horizon, then it may indicate that there is no physical structure around the Cauchy horizon. Of course, this is not classically complete, but by choosing a suitable boundary condition of the quantum gravitational wave function, we can smoothly terminate physical geometries inside the horizon. This choice of boundary condition resolves the mass inflation problem by removing the interior geometry of charged black holes in a quantum-gravitational manner.
@@ VII Conclusion
Analyses of charged black‑hole interiors are inseparable from the instability of the inner Cauchy horizon, where infinite blueshift drives mass inflation and undermines perturbative control. This work advances a complementary quantum‑gravitational resolution: (i) Interpreting the interior Wheeler–DeWitt state as an “annihilation‑to‑nothing” bounce that places both the event and Cauchy horizons in the causal past of the quantum turning point. (ii) Imposing a purely incoming boundary condition at the event horizon then smoothly terminates the interior geometry before a physical Cauchy horizon forms, thus eliminating the mass-inflation region and preventing the associated instabilities. Whether these mechanisms extend to other inner‑horizon spacetimes, including Kerr and regular black holes, remains an open question and a promising avenue for future investigation.
@@
Overall, this framework unifies the classical and quantum descriptions of black-hole interiors, highlights the influence of electromagnetic gauge field on trajectories and localization, and indicates that monotonic decay, quantum bounce, or annihilation is a generic feature of static black holes. Furthermore,the wave function of the Schwarzschild black hole, obtained as the charge-neutral limit of the Reissner–Nordström black hole, exhibits a monotonically decaying behavior and is no longer unbounded. The fact that we can provide three interpretations, where wave packets can be bounded for all cases, is a new observation of this Reissner–Nordström black hole. The choice of boundary condition can resolve the instability issue of the inner horizon, as one can physically remove the inside structure or locate the inner horizon as a causal past. Still, we cannot be sure whether we can generalize this rescue to highly dynamic situations, but we believe that it can open a window to resolving the inner horizon instability issue. For further investigations, we leave the topic for future projects.

### v2

@@ I Introduction
In relation to the Cauchy horizon and the phenomenon of mass inflation, a complementary quantum-gravitational suppression is proposed: (i) Interpret the interior wave function as imposing an “annihilation-to-nothing” boundary condition, so that both the event and Cauchy horizons lie in the causal past of the bounce. In this picture, the infinite blueshift at the would-be Cauchy horizon is not operationally relevant. (ii) With only an ingoing mode from the event horizon, no physical structure forms at the Cauchy horizon. While not a classical completion, an appropriate boundary condition on the wave function smoothly ameliorates the interior geometry and removes the mechanism for mass inflation. Whether this mechanism extends to rotating or regular black holes with inner horizons (e.g., Kerr or regular charged solutions) remains open, and we leave a systematic treatment of these cases to future work.
@@ II Wheeler-DeWitt Equation for Static Black Hole Interiors
where $r_{Q}^{2}={q^{2}}/{4\pi\epsilon_{0}}$ is a characteristic length scale with respect to charge $q$.
The metric ansatz of the spacetime inside a static black hole takes the form
@@ IV.3 Wave Function and Its Boundary Conditions
We note that the choice of the width parameter $\sigma$ for the wave function $\Psi_{-}$ presents a numerical challenge. A small value of $\sigma$ yields a sharply localized Gaussian but can lead to numerical instability, while a large $\sigma$ improves numerical stability at the cost of producing an undesirably broad wave packet. In the regime $r_{Q}\ll r_{s}$, the inner-horizon radius $r_{-}$ becomes extremely small, so that the Gaussian factor $\exp(\sigma^{2}(\lambda-r_{-})^{2}/2)$ is effectively displaced from the intended location. To mitigate this, we adopt a larger $\sigma$ together with a larger $r_{Q}$. This adjustment, however, reduces the decay rate of the wave amplitude in the QQ–direction, as seen in Fig. 3, leading to partial overlap between the two horizon-sourced components. This overlap should not be interpreted as a genuine physical effect of the Reissner–Nordström interior, but rather as a consequence of numerical limitations, see Appendix A; in the ideal physical scenario, the two waves from $r_{\pm}$ would remain sharply localized and well-separated, as depicted in Fig. 2.
@@ IV.4 Neutral Black Hole Limit
In the monotonic‑decay scenario, Schwarzschild and Schwarzschild–(anti)‑de Sitter Wheeler–DeWitt solutions satisfy the DeWitt boundary condition at the curvature singularity, realizing singularity avoidance; however, prior analyses achieve this at the cost of an off‑shell growth that makes the wave function unbounded (see Sec. III.D of Ref. Bouhmadi-Lopez:2019kkt and Sec. III.A of Ref. Chien:2023kqw under the purely incoming horizon condition). In contrast, by reinterpreting the Schwarzschild black hole as the charge‑neutral limit of the Reissner–Nordström black hole and implementing a Gaussian wave with charge fluctuations as in Eq. (33), the neutral‑limit state is normalizable and exhibits strict monotonic decay. This construction furnishes a bounded realization of the monotonic‑decay solution while preserving the DeWitt boundary condition, therefore providing a new perspective on neutralization and refining the late‑time Schwarzschild end‑state picture (cf. Refs. Bouhmadi-Lopez:2019kkt ; Chien:2023kqw ), as illustrated in Fig. 4).
@@ VI Comments on Cauchy Horizon and Mass Inflation
Furthermore, if we think that there is only one piece of the incoming wave from the event horizon, then it may indicate that there is no physical structure around the Cauchy horizon. Of course, this is not classically complete, but by choosing a suitable boundary condition of the quantum gravitational wave function, we can smoothly terminate physical geometries inside the horizon. This choice of boundary condition suppresses the mass inflation problem by removing the interior geometry of charged black holes in a quantum-gravitational manner.
@@ VII Conclusion
Analyses of charged black‑hole interiors are inseparable from the instability of the inner Cauchy horizon, where infinite blueshift drives mass inflation and undermines perturbative control. This work advances a complementary quantum‑gravitational suppression: (i) Interpreting the interior Wheeler–DeWitt state as an “annihilation‑to‑nothing” bounce that places both the event and Cauchy horizons in the causal past of the quantum turning point. (ii) Imposing a purely incoming boundary condition at the event horizon then smoothly terminates the interior geometry before a physical Cauchy horizon forms, thus eliminating the mass-inflation region and preventing the associated instabilities. Whether these mechanisms extend to other inner‑horizon spacetimes, including Kerr and regular black holes, remains an open question and a promising avenue for future investigation.
While the general features summarized in Table 1 for small $\Lambda$ and $r_{Q}$ are expected, it is crucial to underscore that these outcomes are contingent upon the boundary conditions at the inner horizon and the singularity. Acknowledging that the Kantowski-Sachs metric may not fully capture the global properties of the quantized black hole, our observations remain tentative: Schwarzschild and Schwarzschild–(anti-)de Sitter geometries appear to favor monotonic decay, whereas Reissner–Nordström black holes manifest both monotonic decay and ’annihilation-to-nothing’ channels.
A persistent challenge lies in consistently matching these interior results to the regions outside the event horizon and inside the Cauchy horizon, going beyond the specific boundary prescriptions adopted in this study. A robust resolution might involve employing Kruskal-Szekeres coordinates, as demonstrated in Ref. Lin:2025jmz ; however, this approach introduces nontrivial metric parameters, a complexity we reserve for future investigation.
@@
Overall, this framework unifies the classical and quantum descriptions of black-hole interiors, highlights the influence of electromagnetic gauge field on trajectories and localization, and indicates that monotonic decay, quantum bounce, or annihilation is possibly a generic feature of static black holes. Furthermore, the wave function of the Schwarzschild black hole, obtained as the charge-neutral limit of the Reissner–Nordström black hole, exhibits a monotonically decaying behavior and is no longer unbounded. The fact that we can provide three interpretations, where wave packets can be bounded for all cases, is a new observation of this Reissner–Nordström black hole. The choice of boundary condition can suppress the instability issue of the inner horizon, as one can physically remove the inside structure or locate the inner horizon as a causal past. Still, we cannot be sure whether we can generalize this rescue to highly dynamic situations, but we believe that it can open a window to resolving the inner horizon instability issue. For further investigations, we leave the topic for future projects.
Appendix A Numerical Instability of Gaussian Wave
To mitigate the numerical instabilities associated with the Gaussian wavepacket in Eq. (32), we modify the profile function $g(\lambda)$ in Eq. (28) as follows:
$g(\lambda)$ $=\exp\left[-\left(\frac{|\lambda|-\lambda_{0}}{\delta}\right)^{2}\right]\exp\left[-2\left(Y-X+\ln Q-\ln\left(\frac{r_{q}}{r_{s}}\right)\right)^{2}\right].$ (34)
The secondary exponential factor enforces localization along the classical trajectory defined in Eq. (15) for $\Psi_{-}$. As illustrated in Fig. 6, this specific choice constrains the width of the wavepacket $\Psi_{-}$, effectively suppressing numerical overlap. While this approach necessitates an cutoff and careful fine-tuning of the amplitude, it ensures a stable and computationally tractable evolution of the wavepacket.
Figure 6: Numerical integration profiles for the parameter set $\{r_{Q},\sigma,\delta\}=\{0.2,1,0.6\}$, evaluated within the boundary range $(\lambda_{0}+3\delta),\lambda_{0}+3\delta]$. The left panel displays the wavepacket $\Psi_{+}$, while the right panel corresponds to $\Psi_{-}$.